# PhysDepth: Plug-and-Play Physical Refinement for Monocular Depth Estimation in Challenging Environments

### v2

@@ 4.1 Quantitative Results
These physical prior are more effective than other previously adopted domain-specific features in improving MDE (ADDS [liu2021self], and ADFA [vankadari2020unsupervised]).
Among Tables 2, 3, md4All [gasperini2023robust] and Syn2Real-Depth [yan2025synthetic] are two strong baselines, which both adopt GAN-based model and odometry information to achieve better performance.
WSGD [vankadari2023sun] uses three techniques to allow the existing photometric loss to work for both day and challenging condition images.
However, either photometric loss or additional constraints still ignore the light attenuation, leading to suboptimal results. Meanwhile, STEPS [zheng2023steps] learns nighttime condition image enhancer to improve nighttime MDE.
R4Dyn [gasperini2021r4dyn]’s good performance mainly comes from weak radar supervision. However, PhysDepth outperformed R4Dyn because PhysDepth learns robust physical features while also using weak velocity supervision. RNW [wang2021regularizing] is another strong baseline, which performs better than other SOTA by using image enhancement module to enhance image visibility and contrast.
@@ 4.3 Plugin Analysis
Syn2Real-Depth [yan2025synthetic] 0.110 3.455 0.862 0.171 7.689 0.732 0.133 6.926 0.817
PhysDepth + S2R [yan2025synthetic] 0.105 3.450 0.869 0.166 7.681 0.737 0.129 6.921 0.821
MonoDepth [liu2021self] 0.453 11.420 0.700 0.147 6.941 0.827 0.157 7.453 0.795
PhysDepth + MonoDepth [liu2021self] 0.448 11.413 0.708 0.143 6.935 0.823 0.152 7.447 0.799
md4All-AD [gasperini2023robust] 0.124 3.612 0.858 0.182 6.372 0.753 0.141 6.915 0.830
PhysDepth + md4All-AD[gasperini2023robust] 0.119 3.607 0.862 0.179 6.367 0.758 0.137 6.910 0.836
md4All-DD [gasperini2023robust] 0.122 3.604 0.849 0.128 6.607 0.845 0.214 8.376 0.680
PhysDepth + md4All-DD[gasperini2023robust] 0.117 3.600 0.853 0.125 6.601 0.849 0.210 8.371 0.687
DepthDark [zeng2025depthdark] 0.157 4.284 0.760 0.210 7.689 0.630 - - -
PhysDepth+DepthDark [zeng2025depthdark] 0.153 4.280 0.764 0.206 7.683 0.635
@@ 5 Ablation Study
Table 6: Ablation Study - RobotCar-Night [wang2021regularizing]: We evaluate different backbone architectures and module combinations. The best results are in bold for each metric.
@@
Table 7: Ablation Study of Using Different Channel Images on RobotCar-Night Dataset [RobotCarDatasetIJRR]: R means Red Channel, so do B (Blue Channel) and G (Green Channel), RGB. The best results are in bold.
@@
Table 8: Ablation Study of Attenuation Item: Using red channel images from the RobotCar-Night dataset [wang2021regularizing] and the depth range is up to 80m. Linear means that we treat light attenuation as a linear function. And so do Quadratic and Exponential. The best results are in bold.
@@
We also conduct a study of attenuation item $e^{-\mu d_{R}}$ (Exponential) in Table 8. While the other parts in Equation 3 remain the same, we first change $e^{-\mu d_{R}}$ to linear function (Equation 8). And then, we change $e^{-\mu d_{R}}$ to quadratic function (Equation 9). For the experiment settings, we set $a=0.0095,b=0.05$ for Equation 8 [attenuation]. We then set $a=0.0095,b=0.05,c=1.0$ for Equation 9 [attenuation]. All other experiment settings stay the same as in Table 2. The results demonstrate our design choice of using $e^{-\mu d_{R}}$ as attenuation item.
@@ 9 PhysDepth Design
In our base MDE model, the encoder is followed by a self-attention and a series of convolutional upsampling layers to form the decoder Regarding the prediction layers, like Zhou et al. [zhou2017unsupervised], we use $\frac{1}{\alpha\cdot\text{Sigmoid}(x)+\beta}$ with $\alpha=10$ and $\beta=0.0125$ to constrain the predicted depth to be always positive within a reasonable range (e.g., $80$ meters). In addition, we use a Self-Attention module between the encoder and decoder to further improve the accuracy of the baseline.
Encoder: Following follow Piccinelli et al. [piccinelli2024unidepth], for ViT and ConvNeXt we remove all classification-specific layers (e.g., global pooling, fully connected, and softmax) from each. Then we first take the encoder output, flatten it, and apply LayerNorm [ba2016layer]. We then project the feature to a 1024-dimensional channel space using a linear layer and bilinearly interpolate the result to a same resolution of H and W as ViT encoder’s output dimension
Pose Estimation CNN: The pose estimation network learns a pose transformation between two consecutive frames, which serves as a part of our self-supervised signals for depth estimation. We follow the same network in SfMLearner [zhou2017unsupervised], in which we concatenate two consecutive frames along the channel direction. The concatenated result is fed into the Pose Estimation CNN. The outputs are the relative poses between each of the source views and the target view. Specifically, the network consists of 2 convolutions followed by a $1\times 1$ convolution with $6\times 1$ output channels. One channel corresponds to 3 Euler angles and a 3-D translation for each source view. All convolution layers are followed by ReLU except for the last layer, where no nonlinear activation is applied. Since we use photometric reconstruction loss function $\mathcal{L}_{\text{p}}$ from SfMLearner [zhou2017unsupervised] during the training processes, which implies that the rigid assumptions in SfMLearner [zhou2017unsupervised] also apply to our model.
Weak Velocity Supervision: We follow [guizilini20203d, gasperini2023robust] to use a weak velocity supervision $\mathcal{L}_{v}$ for scale awareness and ensure consistent predictions when the photometric consistency does not hold. $\mathcal{L}_{v}$ is defined in Equation 7
@@ 10 DataSet Details
nuScenes-Night and nuScenes-Rain: From the original nuScenes dataset [caesar2020nuscenes], Wang et al. [wang2021regularizing] created the nuScenes-Night dataset split, which contains over 10000 training images and 500 test samples. All images are cropped to $1536\times 768$. All images are in $1242\times 375$.
Table 10: Ablation Study: We test PhysDepth and Plugin-enhanced Models on all images from nuScenes-Night dataset [caesar2020nuscenes] in 40 meters depth range. The best results are in bold.
@@
Table 11: Ablation Study: We test PhysDepth and Plugin-enhanced Models on all images from nuScenes-Night dataset [caesar2020nuscenes] in 50 meters depth range. The best results are in bold.
@@
Table 12: Ablation Study: We test PhysDepth and Plugin-enhanced Models on all images from nuScenes-Night dataset [caesar2020nuscenes] in 60 meters depth range. The best results are in bold.
@@ 11.1 Evaluation over Different Distances
Table 13: Ablation Study: PhysDepth and Plugin-enhanced Models on all images from the RobotCar-Night dataset [wang2021regularizing] and set depth range up to 40 meters. The best results are in bold.
@@
Table 14: Ablation Study: PhysDepth and Plugin-enhanced Models on all images from the RobotCar-Night dataset [wang2021regularizing] and set depth range up to 60 meters. The best results are in bold.
@@
Table 15: Ablation Study: PhysDepth and Plugin-enhanced Models on all images from the RobotCar-Night dataset [wang2021regularizing] and set depth range up to 80 meters. The best results are in bold.
@@
Table 10, 11, and 12 are the results of nuScenes-Night dataset [caesar2020nuscenes] with different depth scales. Specifically, we alter the $\beta$ value from 80 meters to 40, 50, and 60 meters, respectively.
@@
Table 13, 14, and 15 show the results of RobotCar-Night dataset [RobotCarDatasetIJRR] with depth scales of 40, 60, and 80 meters, respectively. The results affirm that PhysDepth can accurately estimate the depth under challenging environments despite the varying depth scales. Notably, in Table 13, we significantly outperform Monodepth2 in RMSE.
@@ 11.2 Additional Ablation Study
Table 16: Ablation Study of $f(R)$ and $R$: Using red channel images from the RobotCar-Night dataset [wang2021regularizing] and the depth range is up to 80m. The best results are in bold.
@@
Table 17: Ablation Study of $\mu$: Using different $\mu$ on RobotCar-Night dataset [wang2021regularizing] and the depth range is up to 80m. The best results are in bold.
@@
Table 18: Ablation Study of $\lambda$: Using different $\lambda$ on RobotCar-Night dataset [wang2021regularizing] and the depth range is up to 80m. The best results are in bold.
@@
We also test PhysDepth with multiple views (3, and 5 views) as input. However, the results exhibit negligible improvements, aligning with the conclusions of Ummenhofer et al. [ummenhofer2017demon] and Zhou et al. [zhou2017unsupervised].
Table 19: Ablation Study of $g$: Using different $g$ on RobotCar-Night dataset [wang2021regularizing] and the depth range is up to 80m. The best results are in bold.
@@
Table 20: Quantitative Results - KITTI [geiger2012we]: Comparison of our model on the KITTI test split by Eigen [eigen2014depth]. M means monocular image, and S means stereo image. The best results are in bold for each metric. All results presented here are without post-processing.
@@
DynaDepth [feng2022disentangling] S $192\times 640$ 0.109 4.705 0.869
MonoDepth2 [godard2019digging] S $320\times 1024$ 0.107 4.764 0.874
EPCDepth [peng2021excavating] S $320\times 1024$ 0.091 4.207 0.901
DepthFormer [li2022depthformer] S $320\times 1024$ 0.090 4.149 0.905
PlaneDepth [wang2023planedepth] S $384\times 1280$ 0.083 3.919 0.913
@@
SfMLearner [zhou2017unsupervised] M $128\times 416$ 0.198 6.565 0.718
DeFeat-Net [spencer2020defeat] M $352\times 480$ 0.126 5.035 0.862
CADepth-Net [yan2021channel] M $192\times 640$ 0.105 4.535 0.892
FM [shu2020feature] M $192\times 640$ 0.104 4.481 0.893
GCP-MonoViT [moon2024ground] M $192\times 640$ 0.096 4.327 0.904
@@
MonoDepth2 [godard2019digging] M $320\times 1024$ 0.115 4.701 0.879
DynamicDepth[feng2022disentangling] M $192\times 640$ 0.103 5.867 0.897
CADepth-Net [yan2021channel] M $320\times 1024$ 0.102 4.407 0.898
DIFFNet [park2021diffnet] M $320\times 1024$ 0.097 4.345 0.907
@@ 12 Additional Qualitative Results
In Figures 7, we use nuScence-Night dataset to compare PhysDepth with SOTA methods: Monodepth2 [godard2019digging] and md4All [gasperini2023robust]. From the results, PhysDepth shows sharper details for the roads and cars, showcasing PhysDepth’s ability to accurately handle varying light conditions under challenging scenarios.
@@ 13 Evaluation on Daytime Dataset
We use KITTI [geiger2012we] dataset to verify the performance of PhysDepth under daytime environments. Table 20 highlights that our model outperforms all SOTA models across all metrics, emphasizing its overall generality and accuracy under daytime environments.
@@
Figure 9: Failure Cases - nuScence-Night: Comparison between PhysDepth and md4All [gasperini2023robust]. Red boxes show the halos because of road lamps and car headlights in the input image, which may cause failure for both PhysDepth and md4All.
@@ 14 Failure Cases
The halo issue is commonly known as one of the most challenging conditions in nighttime depth estimation [wang2021regularizing]. From Figure 9, we can see that SOTA method md4All [gasperini2023robust] produces similar depth results as PhysDepth, demonstrating that SOTA methods may not be able to handle the failure case well.

### v3

@@ 4.1 Quantitative Results
These physical prior are more effective than other previously adopted domain-specific features in improving MDE (ADDS [24], and ADFA [45]).
Among Tables 2, 3, md4All [12] and Syn2Real-Depth [55] are two strong baselines, which both adopt GAN-based model and odometry information to achieve better performance.
WSGD [46] uses three techniques to allow the existing photometric loss to work for both day and challenging condition images.
However, either photometric loss or additional constraints still ignore the light attenuation, leading to suboptimal results. Meanwhile, STEPS [60] learns nighttime condition image enhancer to improve nighttime MDE.
R4Dyn [11]’s good performance mainly comes from weak radar supervision. However, PhysDepth outperformed R4Dyn because PhysDepth learns robust physical features while also using weak velocity supervision. RNW [49] is another strong baseline, which performs better than other SOTA by using image enhancement module to enhance image visibility and contrast.
@@ 4.3 Plugin Analysis
Syn2Real-Depth [55] 0.110 3.455 0.862 0.171 7.689 0.732 0.133 6.926 0.817
PhysDepth + S2R [55] 0.105 3.450 0.869 0.166 7.681 0.737 0.129 6.921 0.821
MonoDepth [24] 0.453 11.420 0.700 0.147 6.941 0.827 0.157 7.453 0.795
PhysDepth + MonoDepth [24] 0.448 11.413 0.708 0.143 6.935 0.823 0.152 7.447 0.799
md4All-AD [12] 0.124 3.612 0.858 0.182 6.372 0.753 0.141 6.915 0.830
PhysDepth + md4All-AD[12] 0.119 3.607 0.862 0.179 6.367 0.758 0.137 6.910 0.836
md4All-DD [12] 0.122 3.604 0.849 0.128 6.607 0.845 0.214 8.376 0.680
PhysDepth + md4All-DD[12] 0.117 3.600 0.853 0.125 6.601 0.849 0.210 8.371 0.687
DepthDark [56] 0.157 4.284 0.760 0.210 7.689 0.630 - - -
PhysDepth+DepthDark [56] 0.153 4.280 0.764 0.206 7.683 0.635
@@ 5 Ablation Study
Table 6: Ablation Study - RobotCar-Night [49]: We evaluate different backbone architectures and module combinations. The best results are in bold for each metric.
@@
Table 7: Ablation Study of Using Different Channel Images on RobotCar-Night Dataset [27]: R means Red Channel, so do B (Blue Channel) and G (Green Channel), RGB. The best results are in bold.
@@
Table 8: Ablation Study of Attenuation Item: Using red channel images from the RobotCar-Night dataset [49] and the depth range is up to 80m. Linear means that we treat light attenuation as a linear function. And so do Quadratic and Exponential. The best results are in bold.
@@
We also conduct a study of attenuation item $e^{-\mu d_{R}}$ (Exponential) in Table 8. While the other parts in Equation 3 remain the same, we first change $e^{-\mu d_{R}}$ to linear function (Equation 8). And then, we change $e^{-\mu d_{R}}$ to quadratic function (Equation 9). For the experiment settings, we set $a=0.0095,b=0.05$ for Equation 8 [20]. We then set $a=0.0095,b=0.05,c=1.0$ for Equation 9 [20]. All other experiment settings stay the same as in Table 2. The results demonstrate our design choice of using $e^{-\mu d_{R}}$ as attenuation item.
@@ 9 PhysDepth Design
In our base MDE model, the encoder is followed by a self-attention and a series of convolutional upsampling layers to form the decoder Regarding the prediction layers, like Zhou et al. [62], we use $\frac{1}{\alpha\cdot\text{Sigmoid}(x)+\beta}$ with $\alpha=10$ and $\beta=0.0125$ to constrain the predicted depth to be always positive within a reasonable range (e.g., $80$ meters). In addition, we use a Self-Attention module between the encoder and decoder to further improve the accuracy of the baseline.
Encoder: Following follow Piccinelli et al. [35], for ViT and ConvNeXt we remove all classification-specific layers (e.g., global pooling, fully connected, and softmax) from each. Then we first take the encoder output, flatten it, and apply LayerNorm [1]. We then project the feature to a 1024-dimensional channel space using a linear layer and bilinearly interpolate the result to a same resolution of H and W as ViT encoder’s output dimension
Pose Estimation CNN: The pose estimation network learns a pose transformation between two consecutive frames, which serves as a part of our self-supervised signals for depth estimation. We follow the same network in SfMLearner [62], in which we concatenate two consecutive frames along the channel direction. The concatenated result is fed into the Pose Estimation CNN. The outputs are the relative poses between each of the source views and the target view. Specifically, the network consists of 2 convolutions followed by a $1\times 1$ convolution with $6\times 1$ output channels. One channel corresponds to 3 Euler angles and a 3-D translation for each source view. All convolution layers are followed by ReLU except for the last layer, where no nonlinear activation is applied. Since we use photometric reconstruction loss function $\mathcal{L}_{\text{p}}$ from SfMLearner [62] during the training processes, which implies that the rigid assumptions in SfMLearner [62] also apply to our model.
Weak Velocity Supervision: We follow [16, 12] to use a weak velocity supervision $\mathcal{L}_{v}$ for scale awareness and ensure consistent predictions when the photometric consistency does not hold. $\mathcal{L}_{v}$ is defined in Equation 7
@@ 10 DataSet Details
nuScenes-Night and nuScenes-Rain: From the original nuScenes dataset [3], Wang et al. [49] created the nuScenes-Night dataset split, which contains over 10000 training images and 500 test samples. All images are cropped to $1536\times 768$. All images are in $1242\times 375$.
Table 10: Ablation Study: We test PhysDepth and Plugin-enhanced Models on all images from nuScenes-Night dataset [3] in 40 meters depth range. The best results are in bold.
@@
Table 11: Ablation Study: We test PhysDepth and Plugin-enhanced Models on all images from nuScenes-Night dataset [3] in 50 meters depth range. The best results are in bold.
@@
Table 12: Ablation Study: We test PhysDepth and Plugin-enhanced Models on all images from nuScenes-Night dataset [3] in 60 meters depth range. The best results are in bold.
@@ 11.1 Evaluation over Different Distances
Table 13: Ablation Study: PhysDepth and Plugin-enhanced Models on all images from the RobotCar-Night dataset [49] and set depth range up to 40 meters. The best results are in bold.
@@
Table 14: Ablation Study: PhysDepth and Plugin-enhanced Models on all images from the RobotCar-Night dataset [49] and set depth range up to 60 meters. The best results are in bold.
@@
Table 15: Ablation Study: PhysDepth and Plugin-enhanced Models on all images from the RobotCar-Night dataset [49] and set depth range up to 80 meters. The best results are in bold.
@@
Table 10, 11, and 12 are the results of nuScenes-Night dataset [3] with different depth scales. Specifically, we alter the $\beta$ value from 80 meters to 40, 50, and 60 meters, respectively.
@@
Table 13, 14, and 15 show the results of RobotCar-Night dataset [27] with depth scales of 40, 60, and 80 meters, respectively. The results affirm that PhysDepth can accurately estimate the depth under challenging environments despite the varying depth scales. Notably, in Table 13, we significantly outperform Monodepth2 in RMSE.
@@ 11.2 Additional Ablation Study
Table 16: Ablation Study of $f(R)$ and $R$: Using red channel images from the RobotCar-Night dataset [49] and the depth range is up to 80m. The best results are in bold.
@@
Table 17: Ablation Study of $\mu$: Using different $\mu$ on RobotCar-Night dataset [49] and the depth range is up to 80m. The best results are in bold.
@@
Table 18: Ablation Study of $\lambda$: Using different $\lambda$ on RobotCar-Night dataset [49] and the depth range is up to 80m. The best results are in bold.
@@
We also test PhysDepth with multiple views (3, and 5 views) as input. However, the results exhibit negligible improvements, aligning with the conclusions of Ummenhofer et al. [44] and Zhou et al. [62].
Table 19: Ablation Study of $g$: Using different $g$ on RobotCar-Night dataset [49] and the depth range is up to 80m. The best results are in bold.
@@
Table 20: Quantitative Results - KITTI [13]: Comparison of our model on the KITTI test split by Eigen [8]. M means monocular image, and S means stereo image. The best results are in bold for each metric. All results presented here are without post-processing.
@@
DynaDepth [10] S $192\times 640$ 0.109 4.705 0.869
MonoDepth2 [15] S $320\times 1024$ 0.107 4.764 0.874
EPCDepth [34] S $320\times 1024$ 0.091 4.207 0.901
DepthFormer [23] S $320\times 1024$ 0.090 4.149 0.905
PlaneDepth [50] S $384\times 1280$ 0.083 3.919 0.913
@@
SfMLearner [62] M $128\times 416$ 0.198 6.565 0.718
DeFeat-Net [41] M $352\times 480$ 0.126 5.035 0.862
CADepth-Net [53] M $192\times 640$ 0.105 4.535 0.892
FM [40] M $192\times 640$ 0.104 4.481 0.893
GCP-MonoViT [28] M $192\times 640$ 0.096 4.327 0.904
@@
MonoDepth2 [15] M $320\times 1024$ 0.115 4.701 0.879
DynamicDepth[10] M $192\times 640$ 0.103 5.867 0.897
CADepth-Net [53] M $320\times 1024$ 0.102 4.407 0.898
DIFFNet [32] M $320\times 1024$ 0.097 4.345 0.907
@@ 12 Additional Qualitative Results
In Figures 7, we use nuScence-Night dataset to compare PhysDepth with SOTA methods: Monodepth2 [15] and md4All [12]. From the results, PhysDepth shows sharper details for the roads and cars, showcasing PhysDepth’s ability to accurately handle varying light conditions under challenging scenarios.
@@ 13 Evaluation on Daytime Dataset
We use KITTI [13] dataset to verify the performance of PhysDepth under daytime environments. Table 20 highlights that our model outperforms all SOTA models across all metrics, emphasizing its overall generality and accuracy under daytime environments.
@@
Figure 9: Failure Cases - nuScence-Night: Comparison between PhysDepth and md4All [12]. Red boxes show the halos because of road lamps and car headlights in the input image, which may cause failure for both PhysDepth and md4All.
@@ 14 Failure Cases
The halo issue is commonly known as one of the most challenging conditions in nighttime depth estimation [49]. From Figure 9, we can see that SOTA method md4All [12] produces similar depth results as PhysDepth, demonstrating that SOTA methods may not be able to handle the failure case well.